a bullet shot straight up is s feet above the ground in t seconds
[[[[where s=f(t)=-16t^2 + 265t]]]]. find the height of the bullet 3 seconds after it is shot

Answer by checkley75(3666)

S=-16*3^2+256*3
S=-16*9+768
S=-144+768
S=624 FEET IS THE HEIGHT AFTER 3 SECONDS.